OpenStudy (anonymous):

if f(x)=9x+1, find f^-1(f(3))

OpenStudy (amistre64):

f(3) = 28; so im guessing its 28 again, but let me chk

OpenStudy (amistre64):

f(3) = 28 f-1 = (x-1)/9 28-1 ----- = 27/9 = 3 9

OpenStudy (anonymous):

it is 28

OpenStudy (amistre64):

... its 3 :)

OpenStudy (amistre64):

i had the right thought, but the wrong concept lol

OpenStudy (anonymous):

3

OpenStudy (anonymous):

27 divided by 9 is 3 yup right